OpenStudy (anonymous):

Express the following Gamma functions and evaluate them using a table of Gamma function

OpenStudy (anonymous):

\[\int\limits_{0}^{1} x^{2} \left( \ln \frac{ 1 }{ x } \right)^{3} dx \] HINT : Put x = \(e^{-u}\)

OpenStudy (anonymous):

Are you able to make the substitution sucessfully?

OpenStudy (anonymous):

No.., i'm not..,

OpenStudy (anonymous):

Note that x=e^(-u) means ln(x)=-u -ln(x)=u ln(1/x)=u Remembe to change your dx as well

OpenStudy (anonymous):

ok.., i'll try

OpenStudy (anonymous):

i would use integration by parts about 3, or 4 times

OpenStudy (anonymous):

actually we might be able to find a reduction formula for this

OpenStudy (anonymous):

i got \[\Gamma (4) = 3 \Gamma (3) = 6 \Gamma (2)= 6 \Gamma (1) = 6 . 1! = 6\]

OpenStudy (anonymous):

You should get \[\frac{\Gamma(4)}{3^{4}}\]

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

i'll try again

OpenStudy (anonymous):

What integral did you get in terms of u?

OpenStudy (anonymous):

First.., \[\int\limits_{0}^{1} x^{2} \left( \ln \frac{ 1 }{ x } \right)^{2} dx \] Let x = exp (-u), so dx = = exp (-u) du Then \[\int\limits_{0}^{1} (\exp(-u))^{2} \left( \ln \frac{ 1 }{ e^{-u} } \right)^{3} (-e^{-u} du)\] \[- \int\limits_{0}^{1} e^{-2u} e^{-u} (\ln e^{3})^3 du = - \int\limits_{0}^{1} e^{-3u} (u)^{3} du\] let y = 3u, so u = y/3, and du = dy/3 Then \[- \int\limits_{0}^{1} \left( \frac{ y }{ 3 } \right)^{3} e^{-y} \frac{ dy }{ 3 }\] \[-\frac{ 1 }{ 3 } \int\limits_{0}^{1} \frac{ y^{3} }{ 27 } e^{-y} dy\] \[-\frac{ 1 }{ 3\times27 } \int\limits_{0}^{1} y^{3} e^{-y} dy\]

OpenStudy (anonymous):

so p -1 = 3.., p = 3+1 \[\Gamma (3+1) = 3 \Gamma (3)\]

OpenStudy (anonymous):

am I wrong ??

OpenStudy (anonymous):

You didn't need the y substitution. You forgot to change your bounds of integration. From a table I found if a > 0 and n > -1 then we have \[\int\limits_{0}^{\infty}x^{n}e^{-ax}dx=\frac{\Gamma(n+1)}{a^{n+1}}\]

OpenStudy (anonymous):

Oh.., my bad..., thank you for helping me ..,